OpenStudy (anonymous):

Given the two functions below, what is the x-intercept of g(x)? f(x) = x + 14 g(x) = f(-x) a. 0,14 b. -14,0 c.0,-14 d.14,0

OpenStudy (nirmalnema):

G(x)=-x-14 g(x)=0 so f(-x)=14 c option is correct

OpenStudy (anonymous):

Yes! Thank you so much

OpenStudy (nirmalnema):

ur welcome...:)

OpenStudy (anonymous):

its wrong. g(x)=-x+14.so the x intercept is (14,0)